John is planning to retire in 15 years. Money can be deposited at 8% compounded quarterly. What quarterly amount must be deposited at the end of each quarter until John retires so that he can make a withdrawal of $ 23,376 semiannually over the first 5 years of his retirement. Assume that John's first withdrawal occurs at the end of six months after his retirement.
Answer #1

Solution: $1825.13
Working: (1+.08/4)^2-1 = 1.0404 - 1 = .0404 = 4.04%
Setting the equivalence relationship at the end of 15 years provides
A (F/A, 2%, 60) = $23,376 (P/A, 4.04%, 10)
A * 114.0515 = $23,376 * 8.0948
A = $189,224.04 / 114.0515
A = $1825.13